Full bibliographic database for special issue of Experimental Astronomy on ISO

A. Coulais

April 30, 2002

References

Abergel2000ExA [1] A. Abergel, M. A. Miville-Deschênes, F. X. Désert, M. Pérault, H. Aussel, and M. Sauvage, The transient behaviour of the long wavelength channel of ISOCAM, Experimental Astronomy, 10:353–368, Août 2000.

Acosta2000ExA [2] J. A. Acosta-Pulido, C. Gabriel, and H. O. Castañeda, Transient effects in ISOPHOT data:

Status of modelling and correction procedures, Experimental Astronomy, 10:333–346, Août 2000.

Altieri2000ExA [3] B. Altieri, L. Metcalfe, J. Blommaert, D. Elbaz, J. . Starck, and H. Aussel, Ultimate

sensitivity with ISOCAM, Experimental Astronomy, 10:291–303, Août 2000.

Biviano2000ExA [4] A. Biviano, M. Sauvage, P. Gallais, O. Boulade, P. Roman, S. Guest, K. Okumura, and S. Ott, The ISOCAM/LW detector dark current behaviour, Experimental Astronomy, 10:255–277, Août 2000.

Blommaert2000ExA [5] J. A. D. L. Blommaert, L. Metcalfe, B. Altieri, A. Biviano, K. Okumura, R. Siebenmorgen, S. Guest, and S. Ott, The isocam responsivity in orbit. standard star photometry, Experimental Astronomy, 10:241–254, Août 2000.

Boulade2000ExA [6] O. Boulade and P. Gallais, The ISOCAM detectors: an overview, Experimental Astronomy,

10:227–239, Août 2000.

Castaneda2000ExA [7] H. O. Castañeda and U. Klaas, Recognition of space weather impact on the isophot detectors,

Experimental Astronomy, 10:369–380, Août 2000.

Claret2000ExA [8] A. Claret, H. Dzitko, J. Engelmann, and J.-L. Starck, Glitch effects in ISOCAM long wave

detector, Experimental Astronomy, 10:305–318, Août 2000.

Doi2000ExA [9] Y. Doi, S. Makiuti, H. Okuda, T. Nakagawa, H. Shibai, M. Kawada, N. Hiromoto, M. Fujiwara, and K. Okumura, Compact two-dimensional array of stressed ge:ga detectors, Experimental Astronomy, 10:393–401, Août 2000.

Dzitko2000ExA [10] H. Dzitko, A. Claret, and J. Engelmann, Cosmic ray effects on the ISOCAM long wave detector,

Experimental Astronomy, 10:279–290, Août 2000.

Engelbracht2000ExA [11] C. W. Engelbracht, E. T. Young, G. H. Rieke, G. R. lis, J. W. Beeman, and E. E. Haller, Observing and calibration strategies for fir imaging with sirtf, Experimental Astronomy, 10:403–413, Août 2000.

Gabriel2000ExA [12] C. Gabriel and J. A. Acosta-Pulido, Deglitching methods by the ISOPHOT interactive analysis

(PIA), Experimental Astronomy, 10:319–331, Août 2000.

Heras2000ExA [13] A. M. Heras, E. Wieprecht, H. Feuchtgruber, F. Lahuis, K. Leech, R. Lorente, P. Morris, A. Salama, and B. Vandenbussche, The ISO-SWS detectors: Performance trends and space radiation effects, Experimental Astronomy, 10:177–197, Août 2000.

Nieminen2000ExA [14] P. Nieminen and J. Sørensen, ISOCAM long wave detector glitches: Data and ray-tracing simula-

tions, Experimental Astronomy, 10:381–391, Août 2000.

Swinyard2000ExA [15] B. Swinyard, P. Clegg, S. Leeks, M. Griffin, T. Lim, and M. Burgdorf, Space operation and performance of doped germanium photo-conducting detectors in the far infrared: Experience from the ISO LWS, Experimental Astronomy, 10:157–176, Août 2000.

Tiphene2000ExA [16] D. Tiphène, D. Rouan, G. Epstein, and P. Le Coupanec, Modelling transient effects in the IR

array of the short wavelength channel of ISOCAM, Experimental Astronomy, 10:347–351, Août 2000.

Valentijn2000ExA [17] E. A. Valentijn and W. F. Thi, ISO’s short wavelength spectrometer - ultimate sensitivity. reducing

the effects of cosmic weather, Experimental Astronomy, 10:215–225, Août 2000.

Wieprecht2000ExA [18] E. Wieprecht, E. Wiezorrek, and L. Haser, Sws-detector glitch effect correction, Experimental

Astronomy, 10:199–214, Août 2000.

Young2000ExA [19] E. T. Young, Summary of ISO detector workshop, Experimental Astronomy, 10:415–418, Août 2000.